# IAP: Improving Continual Learning of Vision-Language Models via Instance-Aware Prompting

### v1

@@ 3.2.2 Class-Distribution Driven Prompting
$\quad E_{i}^{\prime}=$ $\log\varphi(F_{v}(\hat{x});\mu_{i},\Sigma_{i})$ (8)
$\quad=-\frac{1}{2}\left[(F_{v}(\hat{x})-\mu_{i})^{T}\left(\Sigma_{i}\right)^{-1}(F_{v}(\hat{x})-\mu_{i})\right.$
$\quad+d\log 2\pi+\log|\Sigma_{i}|\Big{]},$
$\hat{x}\in\mathbb{R}^{d}$ denote the image input of a test sample, where $d$ represents the dimensionality of the image features. and $\varphi$ denote the probability density function (PDF) corresponding to the learned distribution associated with task $i$. The IKI mechanism computes the PDF values $E^{\prime}_{i}=\varphi(\hat{x})$ for each task $i$ and determines the maximum score as $E^{\prime}_{max}=\text{max}_{i\in[1,\mathcal{T}]}E^{\prime}_{i}$. Subsequently, it applies the sigmoid function $\sigma$ to get the confidence score, yielding $E_{max}=\sigma E^{\prime}_{max}$, which maps $E^{\prime}_{max}$ to the interval [0,1]. This value $E_{max}$ is then employed as the prompting weight within the attention layers of both the vision transformer and the text transformer. The IA-CDDP we proposed is designed with a two-stage architecture, as shown in Figure 5. In the first stage, the module evaluates the confidence scores of the test sample $\hat{x}$ across the $\mathcal{T}$ tasks. For samples exhibiting significantly high or low confidence scores, the module refrains from assigning a weight to the prompt. Instead, it directly utilizes either the incremental learning model or the original CLIP model, depending on the specific bounds:
    Method Aircraft [27] Caltech101 [10] CIFAR100 [22] DTD [6] EuroSAT [14] Flowers [28] Food [2] MNIST [8] OxfordPet [31] StanfordCars [21] SUN397 [44] Average
@@
Table 1: Comparison with SOTA on MCIL benchmark in terms of “Transfer”, “Average”, and “Last” metrics (%). “Ours” denotes our method. We label the best and second methods with bold and underline styles. The presented results are derived from the Order-I, for Order-II results, please refer to the supplementary materials.
$\hat{E}_{max}=\begin{cases}0&\text{for }E_{max}\leq\text{lower bound}\\ 1&\text{for }E_{max}\geq\text{upper bound}\\ E_{max}&\text{into the second stage}\end{cases}\quad\quad.$ (9)
$\quad E^{\prime}_{i,j}=$ $\log\varphi(F_{v}(\hat{x});\mu_{i,j},\Sigma_{i,j})$ (10)
$\quad=-\frac{1}{2}\left[(F_{v}(\hat{x})-\mu_{i,j})^{T}\left(\Sigma_{i,j}\right)^{-1}(F_{v}(\hat{x})-\mu_{i,j})\right.$
$\quad+d\log 2\pi+\log|\Sigma_{i,j}|\Big{]},$
$\hat{E}_{means}=\frac{1}{K}\sum_{k=1}^{K}\sigma E^{\prime}_{[i]},$ (11)
where $E_{[i]}$ represent the top $K$ confidence scores corresponding to the most similar classes within the distribution. Through the application of the second stage of the IA-CDDP strategy, test samples that are distant from the Gaussian distribution of the entire task can still obtain a reasonable confidence score.
Furthermore, for prompts in text encoder, we follow  [36] and leverage a batch-wise confidence score to align with the vision encoder, which can be denoted as follow:
@@
4 Experiments
4.1 Experimental Setting
@@
Dataset. We evaluate our approach within the context of a Multi-Domain Class-Incremental Learning benchmark comprising 11 diverse datasets: Aircraft [27], Caltech101 [10], CIFAR100 [22], DTD [6], EuroSAT [14], Flowers [28], Food [2], MNIST [8], OxfordPet [31], StanfordCars [21] and SUN397 [44]. These datasets collectively encompass 1201 classes, each characterized by distinct distributions. We follow  [51] and conduct our experiment through Order-I and Order-II. Experiments are done in Order-I by default while the experimental details of Order-II is available in the supplementary materials.
Metrics. We evaluate our method’s performance using three metrics proposed by  [51]: “Transfer”, “Last”, and “Avg”. These metrics are tailored to assess distinct aspects of incremental learning, collectively offering a robust evaluation framework. The details about these metrics are included in the supplementary material.
Comparison methods. We compare our proposed IAP method with two categories of SOTA methods in incremental learning. The first category encompasses full-parameter fine-tuning methods, which typically leverage knowledge distillation or rehearsal-based techniques and update all parameters of the model. Notable approaches in this category include Continual-FT, LwF [26], iCaRL [34], LwF-VR [9], WiSE-FT [42], and ZSCL [51]. The second category comprises PEFT incremental learning methods, which adapt models to new distributions by maintaining only a small set of additional parameters. Prominent methods in this category include L2P [41], DualPrompt [40], S-Prompt [39], MoE-Adapter [47] and DIKI [36]. Our proposed IAP method also falls within the PEFT incremental learning benchmark.
Implementation details. To ensure equitable comparisons, we maintain the same backbone as  [51], utilizing CLIP ViT-B/16 as our vision-language model. We employ Stochastic Gradient Descent (SGD) as the optimizer with an initial learning rate of 5.0, paired with a cosine annealing scheduler to dynamically adjust the learning rate during training. The model is trained for 10 epochs across all tasks. Following  [36], we apply prompts to the first 8 transformer layers of the text encoder, with a prompt length of 8. For the image encoder, we adopt the IA-GP strategy and integrate the prompts into all transformer layers during training, with the Gumbel sampling temperature set to 3.0. In the first stage of the IA-CDDP module, we employ an upper bound of 0.8 and a lower bound of 0.2. In the second stage, we set $K=5$ to select the top five most relevant classes. All experiments are performed on an NVIDIA 4090 GPU.
4.2 Experimental Results
Main results. The main results are presented in Table 1. We provide a detailed performance comparison of our proposed method against other SOTA methods within the MCIL benchmark. At the top of the table, we report the zero-shot generalization ability of the CLIP model and its performance under full-parameter fine-tuning as baselines. Our proposed method, denoted as “Ours” in the table, outperforms all SOTA methods for MCIL across three key metrics. Specifically, our method achieves the highest performance on the “Transfer” metric for 4 out of 11 tasks, the “Average” metric for 7 tasks, and the “Last” metric for 6 tasks. Furthermore, our approach establishes new SOTA performance in terms of the overall mean values across all three metrics. While the MoE-Adapter [47] method demonstrates the closest performance to ours, however, our approach achieves this with a significantly smaller number of learnable parameters. For a detailed analysis of the experimental results, please refer to the subsequent sections.
Method Memory Parameters GPU
LwF [26] 211M 32172 MiB
LwF-VR [9] 211M 32236 MiB
ZSCL [51] 211M 26290 MiB
MoE-Adapter [47] 59.8 M 22358 MiB
Ours 2.4 M 19610 MiB
@@ 4.2 Experimental Results
Table 2: Comparison of the use of memory buffers, the scale of trainable parameters, and the GPU burden during the training process.
Method Transfer Average Last
@@
IAP w/o IA-CDDP 68.6 76.6 85.5
IAP w/o IA-GP 69.1 76.5 85.3
IAP 69.2 76.8 85.7
Table 3: The ablation experiments for each module of our proposed method are presented below. Asterisk (*) denotes the original performance reported in the DIKI paper.

### v2

@@ III-B2 Class-Distribution Driven Prompting
$E_{i}^{\prime}$ $=\log\varphi(F_{v}(\hat{x});\mu_{i},\Sigma_{i})$ (7)
$=-\frac{1}{2}\left[(F_{v}(\hat{x})-\mu_{i})^{T}\left(\Sigma_{i}\right)^{-1}(F_{v}(\hat{x})-\mu_{i})\right.$
$+d\log 2\pi+\log|\Sigma_{i}|\Big{]},$
$\hat{x}\in\mathbb{R}^{d}$ denote the image input of a test sample, where $d$ represents the dimensionality of the image features. and $\varphi$ denote the probability density function (PDF) corresponding to the learned distribution associated with task $i$. The IKI mechanism computes the PDF values $E^{\prime}_{i}=\varphi(\hat{x})$ for each task $i$ and determines the maximum score as $E^{\prime}_{max}=\text{max}_{i\in[1,\mathcal{T}]}E^{\prime}_{i}$. Subsequently, it applies the sigmoid function $\sigma$ to obtain the confidence score, yielding $E_{max}=\sigma E^{\prime}_{max}$, which maps $E^{\prime}_{max}$ to the interval [0,1]. This value $E_{max}$ is then employed as the prompting weight within the attention layers of both the vision transformer and the text transformer. The IA-CDDP we proposed is designed with a two-stage architecture, as shown in Figure 5. In the first stage, the module evaluates the confidence scores of the test sample $\hat{x}$ across the $\mathcal{T}$ tasks. For samples exhibiting significantly high or low confidence scores, the module does not assign a weight to the prompt. Instead, it directly utilizes either the incremental learning model or the original CLIP model, depending on the specific bounds:
$\hat{E}_{max}=\begin{cases}0&\text{for }E_{max}\leq\text{lower bound}\\ 1&\text{for }E_{max}\geq\text{upper bound}\\ E_{max}&\text{into the second stage}\end{cases}\quad\quad,$ (8)
$E^{\prime}_{i,j}$ $=\log\varphi(F_{v}(\hat{x});\mu_{i,j},\Sigma_{i,j})$ (9)
$=-\frac{1}{2}\left[(f(\hat{x})-\mu_{i,j})^{T}\left(\Sigma_{i,j}\right)^{-1}(f(\hat{x})-\mu_{i,j})\right.$
$+d\log 2\pi+\log|\Sigma_{i,j}|\Big{]},$
$\hat{E}_{means}=\frac{1}{K}\sum_{k=1}^{K}\sigma E^{\prime}_{[i]},$ (10)
where $E_{[i]}$ represent the top $K$ confidence scores corresponding to the most similar classes within the distribution. Through the second stage of the IA-CDDP strategy, test samples that are distant from the Gaussian distribution of the entire task can still obtain a reasonable confidence score.
TABLE I: Comparison with SOTA on MDCIL benchmark in terms of “Transfer”, “Average”, and “Last” metrics (%). “Ours” denotes our method. We label the best and second methods with bold and underline styles. The presented results are derived from the Order-I.
    Method Aircraft [51] Caltech101 [52] CIFAR100 [53] DTD [54] EuroSAT [49] Flowers [55] Food [56] MNIST [57] OxfordPet [58] Cars [59] SUN397 [50] Average
@@
Furthermore, for prompts in the text encoder, we follow  [9] and leverage a batch-wise confidence score to align with the vision encoder, which can be denoted as follow:
@@
TABLE II: Comparison with SOTA on MDCIL benchmark in terms of “Transfer”, “Average”, and “Last” metrics (%). “Ours” denotes our method. We label the best and second methods with bold and underline styles. The presented results are derived from the Order-II.
    Method StanfordCars [59] Food [56] MNIST [57] OxfordPet [58] Flowers [55] SUN397 [50] Aircraft [51] Caltech101 [52] DTD [54] EuroSAT [49] CIFAR100 [53] Average
CLIP    Zero-shot 64.7 88.5 59.4 89.0 71.0 65.2 24.3 88.4 44.6 54.9 68.2 65.3
   Full Fine-tune 89.6 92.7 99.6 94.7 97.5 81.8 62.0 95.1 79.5 98.9 89.2 89.2
Transfer    Continual-FT 89.5 59.6 57.9 40.0 46.7 11.1 70.0 30.5 26.6 37.7 46.6
   LwF [28] 87.8 58.5 71.9 46.6 57.3 12.8 81.4 34.5 34.5 46.8 53.2
   iCaRL [29] 86.1 51.8 67.6 50.4 57.9 11.0 72.3 31.2 32.7 48.1 50.9
   LwF-VR [60] 88.2 57.0 71.4 50.0 58.0 13.0 82.0 34.4 29.3 47.6 53.1
   WiSE-FT [61] 87.2 57.6 67.0 45.1 54.0 12.9 78.6 35.5 28.4 44.3 51.1
   ZSCL [3] 88.3 57.5 84.7 68.1 64.8 21.1 88.2 45.3 55.2 68.2 64.1
   L2P [38] 70.6 30.7 78.3 42.8 38.3 17.4 75.3 27.4 23.1 20.7 42.5
   DualPrompt [39] 79.9 46.9 85.2 51.3 45.1 9.3 82.7 29.9 42.9 47.2 52.1
   S-Prompts [40] 59.8 46.2 67.7 47.5 43.8 13.5 76.8 31.4 22.6 43.5 45.3
   MoE-Adapter [48] 88.8 59.5 89.1 69.9 64.4 18.1 86.9 43.7 54.6 68.2 64.3
   DIKI [9] 85.8 55.3 89.5 71.1 62.9 23.7 93.6 42.1 43.4 67.9 63.5
   Ours 85.7 59.4 89.1 71.3 62.7 24.4 94.0 43.8 49.0 68.6 64.9
Average    Continual-FT 42.1 70.5 92.2 80.1 54.5 59.1 19.8 78.3 41.0 38.1 42.3 56.2
   LwF [28] 49.0 77.0 92.1 85.9 66.5 67.2 20.9 84.7 44.6 45.5 50.5 62.2
   iCaRL [29] 52.0 75.9 77.4 74.6 58.4 59.3 11.7 79.6 42.1 43.2 51.7 56.9
   LwF-VR [60] 44.9 75.8 91.8 85.3 63.5 67.6 16.9 84.9 44.0 40.6 51.3 60.6
   WiSE-FT [61] 52.6 79.3 91.9 83.9 63.4 65.2 23.3 83.7 45.4 40.0 48.2 61.5
   ZSCL [3] 81.7 91.3 91.1 91.0 82.9 72.5 33.6 89.7 53.3 62.8 69.9 74.5
   L2P [38] 80.1 87.4 86.7 89.6 76.8 59.1 27.7 79.5 33.9 34.6 26.5 62.5
   DualPrompt [39] 78.6 88.4 89.7 91.7 80.0 62.4 23.2 85.0 41.3 51.6 50.7 67.5
   S-Prompts [40] 79.2 86.5 89.5 87.0 78.2 61.5 25.5 83.6 41.9 36.3 47.2 65.1
   MoE-Adapter [48] 84.9 89.9 89.3 91.4 86.2 72.2 33.4 89.4 53.3 61.4 69.9 74.7
   DIKI [9] 84.8 89.0 91.3 93.2 87.8 72.2 34.0 94.5 50.9 53.3 69.6 74.2
   Ours 82.5 89.2 92.3 93.2 88.0 70.4 34.3 94.4 52.4 57.9 70.2 75.1
Last    Continual-FT 24.0 67.3 99.1 87.4 44.3 67.0 29.5 92.3 61.3 81.0 88.1 67.4
   LwF [28] 34.6 69.6 99.3 88.7 61.1 72.5 32.5 88.1 65.6 90.9 87.9 71.9
   iCaRL [29] 46.0 81.5 91.3 82.8 66.5 72.2 16.3 91.6 68.1 83.2 87.8 71.6
   LwF-VR [60] 27.4 61.2 99.4 86.3 60.6 70.7 23.4 88.0 61.3 84.3 88.1 68.2
   WiSE-FT [61] 35.6 76.9 99.5 89.1 62.1 71.8 27.8 90.8 67.0 85.6 87.6 72.2
   ZSCL [3] 78.2 91.1 97.6 92.5 87.4 78.2 45.0 92.3 72.7 96.2 86.3 83.4
   L2P [38] 80.1 89.1 99.1 93.8 96.2 76.5 40.1 86.9 73.5 86.3 84.2 82.3
   DualPrompt [39] 78.6 88.3 99.2 94.1 96.5 76.8 39.8 89.0 71.6 90.7 84.9 82.8
   S-Prompts [40] 79.2 88.1 99.1 94.3 95.8 76.3 39.9 95.5 70.1 97.6 84.4 83.8
   MoE-Adapter [48] 84.1 88.5 94.0 91.8 94.1 77.8 50.4 93.3 77.1 87.7 86.6 84.1
   DIKI [9] 81.8 88.3 99.3 94.7 97.4 76.8 46.4 96.0 74.2 98.0 86.0 85.4
   Ours 82.5 88.6 99.4 94.9 97.7 76.9 46.1 96.1 74.7 98.0 86.6 85.9
TABLE III: Comparison with SOTA on 16-shot MTIL benchmark in terms of “Transfer”, “Average”, and “Last” metrics (%). “Ours” denotes our method. We label the best and second methods with bold and underline styles.
    Method Aircraft [51] Caltech101 [52] CIFAR100 [53] DTD [54] Flowers [55] Food [56] StanfordCars [59] SUN397 [50] Average
CLIP    Zero-shot 24.8 92.9 68.4 43.8 71.4 85.8 65.8 62.6 64.4
   Full Fine-tune 62.0 96.2 89.6 79.5 97.5 92.7 89.6 81.8 86.1
Transfer    ZSCL [3] 87.3 67.7 45.4 67.8 86.6 59.7 63.4 68.3
   L2P [38] 66.7 54.3 30.6 47.3 71.5 54.6 52.4 53.9
   DualPrompt [39] 78.8 64.4 32.0 51.7 77.5 49.4 51.3 57.9
   S-Prompts [40] 70.3 52.7 31.5 54.8 74.0 55.4 50.0 55.5
   DIKI [9] 92.7 68.8 44.1 70.0 86.2 65.1 65.5 70.3
   Ours 93.2 68.9 44.5 71.4 85.5 66.1 65.0 70.9
Average    ZSCL [3] 33.5 90.5 74.7 58.5 79.7 87.7 64.8 64.8 69.3
   L2P [38] 30.2 84.5 70.1 51.9 69.6 77.1 60.0 55.2 62.3
   DualPrompt [39] 36.5 89.5 72.5 52.7 72.3 80.8 56.1 54.2 64.3
   S-Prompts [40] 30.6 86.8 70.0 51.7 74.3 78.5 60.7 53.0 63.2
   DIKI [9] 41.3 95.3 76.5 58.5 82.2 86.4 68.2 66.6 71.9
   Ours 42.5 94.8 77.6 59.3 82.5 86.5 69.6 65.4 72.5
Last    ZSCL [3] 27.7 90.9 74.4 64.7 90.2 89.2 80.6 74.6 74.0
   L2P [38] 30.2 87.1 75.4 64.7 91.9 86.4 76.1 74.7 73.3
   DualPrompt [39] 36.5 91.0 75.1 65.1 92.9 86.2 76.2 74.2 74.7
   S-Prompts [40] 30.6 89.2 75.8 63.8 93.9 86.2 76.7 73.9 73.8
   DIKI [9] 41.3 95.6 79.0 67.3 94.4 86.8 77.6 74.4 77.1
   Ours 42.5 95.8 78.6 68.1 95.3 87.5 79.2 74.8 77.7
IV Experiments
IV-A Experimental Settings
IV-A1 Dataset
@@
We evaluate our approach in the MTIL benchmark, which includes 11 diverse datasets: Aircraft [51], Caltech101 [52], CIFAR100 [53], DTD [54], EuroSAT [49], Flowers [55], Food [56], MNIST [57], OxfordPet [58], StanfordCars [59] and SUN397 [50]. These datasets collectively encompass 1201 classes, each characterized by distinct distributions. The details of datasets are provided in the supplemental materials.
IV-A2 Metrics
We evaluate our method using three metrics proposed by  [3]: “Transfer”, “Last”, and “Avg”. The “Transfer” metric, unconventional in traditional incremental learning approaches, quantifies the forward forgetting of the model’s zero-shot generalization ability across tasks. Specifically, for a given task $i$, it is computed as the average performance across all unseen tasks $i+1,i+2,\cdots,\mathcal{T}$. The “Last” metric assesses the model’s ability to acquire knowledge from new tasks while minimizing catastrophic forgetting of previously encountered tasks, addressing backward forgetting. For task $i$, it is calculated by averaging performance over all seen tasks $i,i-1,\cdots,1$. The “Avg” metric provides a holistic evaluation by integrating both backward and forward forgetting, computed as the mean performance across all tasks $\mathcal{T}$ at each step of the incremental learning process. These metrics are designed to assess the performance of incremental learning methods, collectively offering a robust evaluation framework.
@@ IV-A2 Metrics
TABLE IV: Comparison of the use of memory buffers and the scale of trainable parameters. $\dagger$ represents ours with 8-layer IA-GP strategy.
Method Mem. Param. Trans. Avg. Lst.
iCaRL [29] 211M 50.4 65.7 80.1
ZSCL [3] 211M 68.1 75.4 83.6
MoE [48] $\times$ 59.8M 68.9 76.7 85.0
DIKI [9] $\times$ 1.8M 67.4 75.7 85.1
[…]
Ours $\times$ 2.4M 69.2 76.8 85.7
IV-A3 Comparison Methods
We compare our method with two categories of methods in incremental learning: 1) Full-parameter fine-tuning methods, which typically leverage knowledge distillation or rehearsal-based techniques and update all parameters: Continual-FT, LwF [28], iCaRL [29], LwF-VR [60], WiSE-FT [61], and ZSCL [3]. 2) PEFT methods, which adapt models to new distributions by maintaining only a small set of parameters: L2P [38], DualPrompt [39], S-Prompt [40], MoE-Adapter [48] and DIKI [9]. Our proposed method also falls within PEFT methods.
TABLE V: The ablation experiments for each strategy of our IAP method. Asterisk (*) denotes the original performance reported in the DIKI paper.
Method Tran. Avg. Lst.
@@ IV-A3 Comparison Methods
Ours w/o IA-CDDP 68.6 76.6 85.5
Ours w/o IA-GP 69.1 76.5 85.3
Ours 69.2 76.8 85.7
IV-A4 Implementation Details
To ensure fair comparisons, we maintain the same backbone as  [3], utilizing CLIP ViT-B/16 as our vision-language model. We employ Stochastic Gradient Descent (SGD) as the optimizer and the model is trained for 10 epochs across all tasks. Following  [9], we set the learning rate to 5.0 and apply prompts to the first 8 transformer layers of the text encoder, with a prompt length of 8. For the image encoder, we adopt the IA-GP strategy and integrate the prompts into all transformer layers, with the Gumbel sampling temperature set to 3.0. The batch size was set to 128 during training and 256 during inference. Experiments ere conducted using FP16 precision to minimize the computational costs of training and inference. In the first stage of the IA-CDDP strategy, we employ an upper bound of 0.8 and a lower bound of 0.2, a perturbation of $1\times 10^{-6}$ is introduced to all sampled values during Gumbel sampling to mitigate numerical instability. In the second stage, we set $K=5$ to select the top five most relevant classes, we also follow  [9] and apply a regularization term of $10^{-7}$ to the task-wise covariance matrix and $10^{-3}$ to the class-wise covariance matrix to address numerical instability. All experiments are performed on an NVIDIA 4090 GPU.
TABLE VI: Comparison of our method with two additional adaptive prompting strategies.
Method Tran. Avg. Lst.
Random Layer 68.8 75.9 84.5
Soft Gumbel 68.6 76.3 85.3
Ours 69.2 76.8 85.7
IV-B Experimental Results
IV-B1 Main Results
We follow  [3] and conduct our main experiment through Order-I task sequence, where the model is trained incrementally on a series of tasks in the following order: Aircraft [51], Caltech101 [52], CIFAR100 [53], DTD [54], EuroSAT [49], Flowers [55], Food [56], MNIST [57], OxfordPet [58], StanfordCars [59] and SUN397 [50]. The results are presented in Table I. We provide a detailed performance comparison of our proposed method against other SOTA methods within the MTIL benchmark. At the top of the table, we report the zero-shot generalization ability of the CLIP model and its performance under full-parameter fine-tuning as baselines. Our proposed method, denoted as “Ours” in the table, outperforms all SOTA methods for MTIL across three key metrics. Specifically, our method achieves the highest performance on the “Transfer” metric for 4 out of 11 tasks, the “Average” metric for 7 tasks, and the “Last” metric for 6 tasks. Furthermore, our approach establishes new SOTA performance in terms of the overall mean values across all three metrics. While the MoE-Adapter [48] method demonstrates the closest performance to ours, our approach achieves this with a significantly smaller number of learnable parameters. More detailed results for each incremental learning session can be found in the supplemental materials.
IV-B2 Order-II Result
We evaluate our method in the second task order, where the tasks are acquired through the following sequence: StanfordCars [59], Food [56], MNIST [57], OxfordPet [58], Flowers [55], SUN397 [50], Aircraft [51], Caltech101 [52], DTD [54], EuroSAT [49] and CIFAR100 [53]. The experimental results are provided in Table II. In the second order, our methods also achieve the optimal performance among all the metrics “Transfer”, “Average” and “Last”, which demonstrates that our method is robust to variations in task order.
IV-B3 Few-shot Incremental Learning
We follow [9] and evaluate our method in the 16-shot MTIL setting, where only 16 samples per task are available during model training. The comparison results are provided in Table III, which uses the same metrics as the main results. Our method also outperforms all other SOTA approaches. This reveals that our method can effectively learn new tasks with limited data while exhibiting minimal forgetting of previously acquired knowledge.
IV-B4 Trainable Parameters and Memory Buffer
[…]
IV-B5 Instance-Aware Gated Baselines
To provide a more intuitive demonstration of the effectiveness of our IA-CP strategy, we conducted two additional baseline experiments with adaptive prompting strategies: 1) Random prompting layers. For each instance, the model randomly determines the number of prompting layers. Specifically, for every transformer layer, there is a 50% probability that a prompt will be injected. 2) Soft Gumbel Softmax: rather than enforcing a binary decision (0 or 1) on whether to inject a prompt into a transformer layer, the model outputs a soft value to control the prompting process. We compared these two baselines against the results of our main experiment, as shown in Figure VI. Under identical experimental settings, our proposed method consistently outperformed both baselines across all the “Transfer” “Average” and “Last” metrics.
(a) fig 5
(b) fig 6
(c) fig 7
(d) fig 1
(e) fig 2
(f) fig 3